# Relative pose of three calibrated and partially calibrated cameras from four points using virtual correspondences

### v2

@@ 3.1 Calibrated cameras
4p3v(M±$\pm\delta$) solver:  While the mean point correspondence used in the 4p3v(M) solver can provide a good approximation of a correct correspondence, such a correspondence can be noisy. In the 4p3v(M±$\pm\delta$) solver, we thus, in addition to the mean point $\mathbf{m}^{2}=\left[x,y\right]$ of three points $\left\{\mathbf{x}_{i}^{2},\mathbf{x}_{j}^{2},\mathbf{x}_{k}^{2}\right\}$ in the second image, generate two additional points. These points are (1) $\mathbf{m}^{2}_{\delta}=\left[x\pm\delta,y\right]$ if the longest dimension of the triangle $\mathcal{T}^{2}=\Delta\left\{\mathbf{x}_{i}^{2},\mathbf{x}_{j}^{2},\mathbf{x}_{k}^{2}\right\}$ is in the x direction or (2) $\mathbf{m}^{2}_{\delta}=\left[x,y\pm\delta\right]$ if it is in the y direction. All three points, i.e., $\mathbf{m}^{2}$ and $\mathbf{m}^{2}_{\delta}$ are placed in correspondence with the mean point $\mathbf{m}^{1}$. The 4p3v(M±$\pm\delta$) solver in the first step calls the 5pt solver [38] three times, with the $5^{th}$ correspondence being either $\mathbf{m}^{1}\leftrightarrow\mathbf{m}^{2}$ or $\mathbf{m}^{1}\leftrightarrow\mathbf{m}^{2}_{\delta}$. The results of these three 5pt solvers are collected to create hypotheses for the relative pose of the first two cameras inside RANSAC. The shift $\delta$ is selected relative to the size of the triangle $\mathcal{T}^{2}$.

### v3

@@ 3.1 Calibrated cameras
4p3v(M±$\pm\delta$) solver:  While the mean point correspondence used in the 4p3v(M) solver can provide a good approximation of a correct correspondence, as mentioned, it can also be noisy. In the 4p3v(M±$\pm\delta$) solver, we thus, in addition to the mean point $\mathbf{m}^{2}=\left[x,y\right]$ of three points $\left\{\mathbf{x}_{i}^{2},\mathbf{x}_{j}^{2},\mathbf{x}_{k}^{2}\right\}$ in the second image, generate two additional points. These points are (1) $\mathbf{m}^{2}_{\pm\delta}=\left[x\pm\delta,y\right]$ if the longest dimension of the triangle $\mathcal{T}^{2}=\Delta\left\{\mathbf{x}_{i}^{2},\mathbf{x}_{j}^{2},\mathbf{x}_{k}^{2}\right\}$ is in the x-direction or (2) $\mathbf{m}^{2}_{\pm\delta}=\left[x,y\pm\delta\right]$ if it is in the y-direction. All three points, i.e., $\mathbf{m}^{2}$, $\mathbf{m}^{2}_{-\delta}$, and $\mathbf{m}^{2}_{+\delta}$ are placed in correspondence with the mean point $\mathbf{m}^{1}$. The 4p3v(M±$\pm\delta$) solver in the first step calls the 5pt solver [34] three times, with the $5^{th}$ correspondence being either $\mathbf{m}^{1}\leftrightarrow\mathbf{m}^{2}$, $\mathbf{m}^{1}\leftrightarrow\mathbf{m}^{2}_{-\delta}$, or $\mathbf{m}^{1}\leftrightarrow\mathbf{m}^{2}_{+\delta}$. The results of these three 5pt solvers are collected to create hypotheses for the relative pose of the first two cameras inside RANSAC. The shift $\delta$ is selected relative to the size of the triangle $\mathcal{T}^{2}$.
$\mathbf{4^{th}}$ point in the third view:  The 4p3v(M) and 4p3v(M±$\pm\delta$) solvers are actually solving the configuration ${(S_{4},S_{4},S_{4}^{3})}$, i.e., they are not using information from the $4^{th}$ point in third view, i.e., point $\mathbf{x}^{3}_{4}$. The information from $\mathbf{x}^{3}_{4}$ can be used in the solver for the ${(S_{4},S_{4},S_{4})}$ configuration in two different ways: Filtering (+F): $\mathbf{x}^{3}_{4}$ can be used to filter out geometrically infeasible solutions returned by the 5pt+P3P solver that is used inside M-based solvers. The 5pt+P3P solver returns multiple solutions that can be evaluated w.r.t. $\mathbf{x}^{3}_{4}$. Since the returned solutions can be affected by larger noise in the mean-point correspondence, we do not simply select the solution with the smallest error on $\mathbf{x}^{3}_{4}$, but we keep all solutions that have an epipolar error on $\mathbf{x}^{3}_{4}$ smaller than twice the threshold used inside RANSAC. Our experiments show that this filtering can improve the speed of the proposed solvers. However, as a trade-off, there is a small drop in the precision of the solvers since, in some cases, geometrically correct solutions are filtered. Refinement (+R): $\mathbf{x}^{3}_{4}$ can be used to refine the solutions returned by the 5pt+P3P solver used inside M-based solvers. These solutions have zero error on the original four points and the mean-point correspondence in the first two views, but can have a large error on $\mathbf{x}^{3}_{4}$. Rather, we want solutions that minimize the epipolar errors on the original four points in the three views. Note that ${(S_{4},S_{4},S_{4})}$ is an overconstrained configuration and thus for noisy data, there is, in general, no solution with zero error on all four points in the three views. We perform refinement of the poses by minimizing the epipolar errors of all original four points in three views using Levengerg-Marquardt optimization (LM), initialized using the solutions from the M-point solvers. Experiments with different numbers of iterations show that two iterations are usually sufficient to obtain an improvement (see SM).